40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11174     Entry time: Wed Mar 25 21:44:20 2015
Author: Koji 
Type: Update 
Category: LSC 
Subject: IFO recovery / PRFPMI locking activity 

[Koji, Den]

- Aligned the arms with ASS. It had alot of offset accumulated. We offloaded it to the suspension.

- We could lock the PRMIsb with the new setup.
PRCL: REFL165I (-14deg, analog +9dB)) -0.1, Locking FM4/5, Triggered FM 2
MICH: REFL165Q (-14deg, analog +9dB) -1.5, Locking FM4/5, Triggered FM2/6/9

- Demod phases at REFL were adjusted such that PRCL in Q signals were minimized :
REFL165 -80deg => -14deg
POP55 -63deg
REFL11 +164 => +7
REFL33 +136 => +133

Note: analog gains: REFL11: +18dB,  REFL33: +30dB, POP55: +12dB, REFL165: +9dB

- Try some transition between REFL signals to check the signal quality.
Measure TFs between the REFL signals

PRCL gain
REFL11I/REFL165I = +58
REFL33I/REFL165I = +8.5
POP55I /REFL165I = -246

MICH gain
REFL11Q/REFL165Q = +11
REFL33Q/REFL165Q = -1.5
POP55Q /REFL165Q = +280

- This resulted us to figure out the relationships of the numbers in the input matrix 

REFL55I/Q -4e-3/4e-3
REFL165I/Q 1.0/1.0 (reference)
REFL11I/Q  0.02/0.1
REFL33I/Q +0.12/-0.7


Full locking trial

Arm locked -> ALS -> Arm offset locked
PRMI locking
REFL165 phase tuned -110deg
PRCL gain -0.1 / MICH gain -2

We needed script editing.
Previous script saved in: /opt/rtcds/caltech/c1/scripts/PRFPMI/carm_cm_up_BACKUP.sh

Change:
- PRMI gain setting (input matrix & servo gain)
- CARM/DARM transition setting (see below)


The current CARM/DARM transition procedure:

== CARM TRANSITION (PART1) ==
- CM REFL1 gain is set to be -32
- CARM_B is engaged and the gain is ramped from 0 to +2.5
- Turn on FM7 (integrator)
- MC IN2 (AO path) engaged
- MC IN2 gain increased from -32 to -21

== DARM TRANSITION (PART1) ==
- DARM_B is engaged and the gain is ramped from 0 to +0.1
- Turn on FM7 (integrator)

== CARM TRANSITION (PART2) ==
- CM REFL1 Gain is increased from -32 to -18
- Ramp down CARM A gain to 0

== DARM TRANSITION (PART2) ==
- DARM_B gain is incrased to 0.37. At the same time DARM_A gain is reduced to 0


We succeeded to make the transition several times in the new setting.

- But later the transition got hard. We started to see big jump of the arm trans (TRX/Y 50->100) at the CARM transition.

- We tested the PRCL transition from 165MHz to 55MHz. 55MHz (i.e. POP55 which is REFL55PD) looks alot better now.

- ~1:30 The PMC was realigned. This  increased PMC_TRANS about 10%. This let the Y arm trans recover ~1.00 for the single arm locking

- Decided to end around 3:00AM

ELOG V3.1.3-